OpenStudy (anonymous):

Suppose a circular platform is rotating at a constant rate of 21.0 revolutions/min. How far away from the centre would one need to stand in order to have a radial acceleration of 9.7 m/s2?

OpenStudy (anonymous):

What is the mathematical relationship that connects rotational motion to radial acceleration?

OpenStudy (michele_laino):

the radial acceleration \(a_r\) has the subsequent magnitude: \[\Large {a_r} = {\omega ^2}r\] where \(r\) is the requested distance, and \(\omega\) is given by the subsequent computation: \[\Large \omega = \frac{{2\pi n}}{{60}} \cong \frac{{6.28 \cdot 21}}{{60}} = ...?\] Of course \(a_r=9.7\;m/s^2\)